A farmer has the opportunity to purchase an advanced guidance system for precision farming. This is the advertisement:

New Outback MAX Guidance System

Outback MAX™ with Outback ConnX™ — introducing the next generation Outback Guidance® system that redefines simplicity in precision farming.

The farmer believes that the guidance system that can be purchased for $5,500 will improve corn production by 220 bushels per year. He plans to keep the Guidance system for 8 years. At the end of 8years he believes he can sell it for $1,000. The farm’s accountant has done some preliminary work. The marginal tax rate is 25%, the pre-tax risk adjusted discount rate is 8% and inflation is assumed to be zero. The account calculates the present value of the tax savings from depreciation to be $1,067.31, and the present value of the after-tax terminal value to be $470.56. The price of corn is expected to be $4.30/bushel over the 8 years.

(i)        Calculate the additional after-tax net returns from corn production if the farmer purchases the guidance system.

A.         $900

B.       $709

C.         $270

D.         $750

E.         None of these answers

Enter Response Here:

(ii)       Calculate the after-tax risk adjusted discount rate.

A. 6.0%

B.         12%

C.         3.6%

D.         1.25%

E.         None of these answers

Enter Response Here:

(iii)      Calculate the present value of the additional after-tax net returns from corn production if the farmer purchases the guidance system.

A.         $5,932

B.         $1,779

*C.       $4,406

D.         $2,963

E.         None of these answers

Enter Response Here:

(iv)      Calculate the NPV of the guidance system.

A.         $156

B.         $1100

C.         $445

D.       $444

E.         None of these answers

Enter Response Here:

(v)       Calculate the break-even improvement in corn production (yield) if the farmer purchases the guidance system.

A.         $150

B.       $198

C.         $140

D.         $170

E.         None of these answers

Please help by explaining the process.
Answer #1

Answer (i):

Correct answer is:

B.  $709

Explanation:

The farmer believes that the guidance system will improve corn production by 220 bushels per year. The marginal tax rate is 25%.

The price of corn is expected to be $4.30/bushel over the 8 years.

Hence:

After-tax net returns from corn production if the farmer purchases the guidance system = (220 * $4.30) * (1 - 25%)

= $946 * 75%

= $709.50

Closest option is B and other options A, C and D are incorrect.

Answer (ii):

Correct answer is:

A. 6.0%

Explanation:

Given:

The marginal tax rate is 25%, the pre-tax risk adjusted discount rate is 8%.

Hence:

After-tax risk adjusted discount rate = 8% * (1 - 25%) = 6%

Option A is correct and other options B, C and D are incorrect.

Answer (iii):

Correct answer is:

C. $4,406

Explanation:

Present value of the additional after-tax net returns = Annual after-tax net returns from corn production * PV of $1 annuity for 8 years at 6% rate

= 709.50 * (1 - 1/ (1 + 6%) ^8)/ 6%

= $4,405.85

or $4,406

Option C is correct and other options A, B and D are incorrect.

Answer (iv):

Correct answer is:

D. $444

Explanation:

Given:

Initial investment = $5,500

Present value of the tax savings from depreciation = $1,067.31

Present value of the after-tax terminal value = $470.56

We calculated in answer iii above:

Present value of the additional after-tax net returns = $4,405.85

Hence:

NPV = $4,405.85 + $1,067.31 + $470.56 - $5,500

= $443.72

or $444

Option D is correct and other options A, B and C are incorrect.

Answer (v):

Correct answer is:

B. 198

Explanation:

At break-even improvement in corn production (yield), NPV will be zero.

Present value of after-tax net returns from corn production will be = $4,405.85 - $443.72 = $3962.13

After tax annual improvement in corn production in dollars = 3962.13 / ((1 - 1/ (1+6%) ^8)/6%) = $638.05

Before tax annual improvement in corn production in dollars = 638.05 / (1 - 25%) = $850.73

Break-even improvement in corn production (yield) in units = 850.73 / 4.30 = 198 Bushels

Option B is correct and other options A, C and D are incorrect.